LABORATORY AND ASTRONOMICAL DETECTION OF CYANIC ACID, HOCN

\underline{SANDRA BRÜNKEN}, CARL A. GOTTLIEB, MICHAEL C. MCCARTHY, PATRICK THAD-DEUS, Harvard Smithsonian Center for Astrophysics, and School of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138, USA; ARNAUD BELLOCHE, KARL M. MENTEN, Max-Planck-Institut für Radioastronomie, 53121 Bonn, Germany.

Isocyanic acid (HNCO) was among the first interstellar molecules identified by radio astronomical observations^{a} and has since been found in at least sixty Galactic sources^{b}, and in three external galaxies^{c}. It is the most stable of four isomers with this elemental composition, and has been extensively studied from the microwave to the IR region, as has the higher lying isomer fulminic acid, HCNO^{d}. Here we report the recent laboratory and astronomical detection of the second-most stable isomer, cyanic acid HOCN, which is only 21 kcal/mol or 0.9 eV higher in energy than HNCO. In the laboratory, 45 a-type transitions were measured in the frequency range 20-360 GHz up to J = 17 and K_{a} = 4 with a Fourier Transform Microwave spectrometer (FTM) or a free-space mm-wave absorption spectrometer. HOCN was produced in a discharge through water and cyanogen. The identification was confirmed by observing lines of DOCN at the expected isotopic shift, and by resolving the ^{14}N hyperfine structure. Based on this laboratory study, we have now detected HOCN towards SgrB2(M) in 3-mm and 2-mm survey data^{e} via four a-type transitions in the K_{a} = 0 and three in the K_{a} = 1 ladder.

aD. Buhl, L. E. Snyder, P. R. Schwartz, and J. Edrich, Nature 243, 513 (1973)

be.g., J. M. Jackson, J. T. Armstrong, and A. H. Barrett, ApJ 280, 608 (1984), I. Zinchenko, C. Henkel, R. Q. Mao, A&A 361, 1079 (2000)

cNguyen-Q-Rieu, C. Henkel, J. M. Jackson, and R. Mauersberger, A&A 241, L33 (1991)

dB. P. Winnewisser, ”Molecular Spectroscopy: Modern Research” (K. Narahari Rao, Ed.), Vol. III, Academic Press, New York (1985)

eA. Belloche, K. M. Menten, private communication; A. J. Remijan, D. P. Leigh, A. J. Markwick-Kemper, B. E. Turner, ArXiv e-prints 0802.2273 (2008)